Find molarity, molality, normality, ppm, mass%, mole fraction in 98% H2SO4 solution. (p=1.8g/ml). Please give the solution.

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello
Assuming you have 98g h2so4 in 100g of solution.
100g => 1*100/1.8ml =55.55ml
molecular weight ofh2so4=98g
=>molarity = 1*1000/55.55 =1.8 molar
=>molality = 1/(55.55*1.8-1*98)/1000=1.10molal
=>normality =98/(98/2)/55.55/1000=1000*2/55.55=36.00
now please other things yourself.
Just try by dealing mass of solute and solvent seperately.